an envelope measures 24 centimeters long and 10 centimeters wide. what is the perimeter of the envelope
Answer by stanbon(75887)

an envelope measures 24 centimeters long and 10 centimeters wide. what is the perimeter of the envelope
---------------------
Perimeter = 2(L + W)
-----------
P = 2(24+10)
P = 2(34)
P = 68 cm
======================
Cheers,
Stan H.